OpenStudy (anonymous):

maggies mold has a right cylindrical shape. the mold is 8 inches high, and the bottom of the mold has a diameter of 5 inches. how many cubic inches are in maggies mold?

OpenStudy (anonymous):

V=h(pi)r^2

OpenStudy (anonymous):

V=8(pi)6.25 V=157.08

Directrix (directrix):

Volume = π r ² h V = π (5/2) ² (8) V = 50 π = 157 approximately

Directrix (directrix):

Questions?

OpenStudy (anonymous):

no